# Random walks and the ”Euclidean” association scheme in finite vector spaces

### 09 Jan 2024 v1

@@ Abstract.
Keywords: Association schemes, Kloosterman sums, Random Walks. 2020 Mathematics Subject Classification: Primary: 05E30, 05C90. Secondary: 11L05, 11T23.
@@ 1. Introduction
In this paper we study the unit random walk in finite vector spaces. We also provide a discussion of a related association scheme that was the subject of the first author’s thesis work [3]. The books [4], [6] are good references for the spectral graph theory concepts used in this paper. The book [4] is a good reference for association schemes. Let $\mathbb{F}_{q}$ be a finite field of odd prime power order $q$ and $V=\mathbb{F}_{q}^{d}$ be the standard $d$-dimensional vector space over $\mathbb{F}_{q}$. Note that all finite vector spaces of odd characteristic are of this form. Equip $V$ with the standard bilinear dot product associated to the nondegenerate quadratic form $Q(v)=v_{1}^{2}+\dots+v_{d}^{2}$. The ”distance” between two points $v$ and $w$ in $V$ as determined by this quadratic form is given by $Q(v-w)$ which is the standard ”Euclidean” distance formula without the square root. The quadratic space $(V,Q)$ will be refered to as $d$-dimensional Euclidean space over $\mathbb{F}_{q}$ in this paper. Note that it is known that there are only two nondegenerate quadratic forms up to isometry on any finite vector space ([14]), besides the one derived from the standard dot product that we are using, the other one is a ”Lorentzian” quadratic form with $Q_{Lorentz}(v)=v_{1}^{2}+\dots+v_{d-1}^{2}+\xi v_{d}^{2}$ where $\xi$ is any fixed nonsquare element in $\mathbb{F}_{q}$. Most of what we discuss will also hold for this other quadratic form but we will stick to the dot product in this paper for brevity. For every $t\in\mathbb{F}_{q}$, we may define the distance-t graph whose vertex set is $V$ and where there is an edge between $v,w\in V$ if and only if $Q(v-w)=t$ and $v\neq w$. This regular graph has a $q^{d}\times q^{d}$ adjacency matrix $\mathbb{A}_{t}$ with respect to some fixed ordering of the vertices of $V$ and common vertex degree $|S_{t}|$ where $S_{t}=\{v\in V\setminus\{0\}|Q(v)=t\}$ is the ”sphere of radius $t$” and has order $q^{d-1}(1+o(1))$ when $t\neq 0$. (The exact order is known but here for brevity we collect secondary terms in $o(1)$ which tends to 00 as $q\to\infty$). In fact this graph is a Cayley graph with connection set $S_{t}$ and hence is in fact vertex-transitive (i.e., there is a graph automorphism taking any vertex to any other vertex). The corresponding Markov chain (see [11] for basic Markov chain terminology) on this graph is called the distance-t random walk in $V$ with the case $t=1$ called the unit random walk in $V$. It corresponds to a situation where at each step, the current state evolves by taking a distance $t$ step with each such step equally likely. The transition matrix for this Markov chain is $T_{t}=\frac{1}{|S_{t}|}\mathbb{A}_{t}$. As explained in more detail within this paper, we show that the matrices $\{\mathbb{A}_{t}|t\in\mathbb{F}_{q}\}$ determine a symmetric association scheme with corresponding real and complex Bose-Messner algebras. What this means is that we have identities of the form
@@
where the intersection numbers $p_{i,j}^{k}$ correspond to the number of ways that a given ”line segment” of length $k$ can be completed to a triangle of side lengths $i,j$ and $k$. We call this particular scheme the Euclidean association scheme for the finite field $\mathbb{F}_{q}$ and compute certain change of basis matrices $P$ and $Q$ which are generally important in the theory of such schemes. We also compute this association scheme’s intersection numbers. This first part of the paper’s results appear in the first author’s thesis and we publish them here as a step towards future results using this association scheme machinery (we only provide one application in this paper for brevity). The eigenvalues of $\mathbb{A}_{t}$ are related to twisted Kloosterman sums:
@@
A second result of this paper calculates $R_{q,\ell,t}$ (see Theorem 7.1), the probability that you return to the vertex you started from after $\ell$-steps in the random distance $t$-walk, in terms of these moments of Kloosterman sums. We show that this probability is independent of your starting state/vertex and is given by:
@@
Note the first $\frac{1}{q^{2}}$ term in these asymptotic formulas is what one would expect if the location after $\ell$ steps were equally likely to be anywhere in the plane and the second term represents an arithmetic bias against that happening. In particular we see that for $\ell\geq 5$, we have $\mathbb{R}_{q,\ell}$ is $\frac{1}{q^{2}}(1+o(1))$ but for smaller $\ell$ arithmetic bias is significant and so we need at least $5$ steps to achieve uniformity. Note $R_{q,1}=0,R_{q,2}=\frac{1}{q}(1+o(1)),R_{q,4}=\frac{3}{q^{2}}(1+o(1))$ are all ”arithmetically biased”. Similar results hold for $q=1\text{ mod }4$ and can be seen in the paper.
@@ 3. Eigenvalues of the Distance $t$-graphs
For any $t\in\mathbb{F}_{q}$, let $X(q,t)$ denote the associated distance $t$-graph on $\mathbb{F}_{q}^{d}$. In this section, we describe the spectrum of the adjacency matrix of this graph, $A_{t}$, when $t\neq\bar{0}$, i.e., for $t\in\mathbb{F}_{q}$. This spectrum was previously computed by various authors for example in [8] but we include a brief discussion here to fit it into the language of association schemes. Standard spectral theory of finite Cayley graphs over Abelian groups, shows that the complex eigenfunctions of this graph’s adjacency operator $A_{t}$ are the characters $\chi_{m}(x)=\chi(m\cdot x)$ as $m$ ranges over $\mathbb{F}_{q}^{d}$. Here $\cdot$ stands for dot product. The corresponding eigenvalue of $A_{t}$ on $\chi_{m}$ which we will denote $\lambda_{m,t,d}$ (recall $d$ is the dimension of $V$) is given by
@@ 8. Intersection Numbers and Matrices of the scheme
We will start with the planar, $d=2$ case. This case was partially computed in [2] in a different context.

### 12 Jan 2024 v2

@@ Abstract.
In this paper, we provide an application to the random distance-$t$ walk in finite planes and derive asymptotic formulas (as $q\to\infty$) for the probability of return to start point after $\ell$ steps based on the ”vertical” equidistribution of Kloosterman sums established by N. Katz. This work relies on a ”Euclidean” association scheme studied in prior work of W.M.Kwok, E. Bannai, O. Shimabukuro and H. Tanaka. We also provide a self-contained computation of the P-matrix and intersection numbers of this scheme for convenience in our application as well as a more explicit form for the intersection numbers in the planar case. Keywords: Association schemes, Kloosterman sums, Random Walks. 2020 Mathematics Subject Classification: Primary: 05E30, 05C90. Secondary: 11L05, 11T23.
@@ 1. Introduction
In this paper we study the unit random walk in finite vector spaces. We also provide a discussion of a related association scheme that was the subject of the first author’s thesis work [4] that we latter were informed was previously studied by W.M. Kwok, E. Bannai, O. Shimabukuro and H. Tanaka in [2], [12]. In [12], the P-matrix and intersection numbers of this scheme were worked out in terms of the character table of the ”Euclidean” group over finite fields but had some errors which were latter corrected in [2] where a connection to Kloosterman sums was also mentioned, including Kloosterman’s original bound. In this paper, we recap these calculations first to connect them more explicitly to our applications which use some deeper distributional data about Kloosterman sums provided by equidistribution results from number theory. (Along the way we also provide an explicit calculation of the planar intersection numbers implicitly described in [12].) The books [5], [7], [17] are good references for the spectral graph theory concepts used in this paper. The book [5] is a good reference for association schemes. Let $\mathbb{F}_{q}$ be a finite field of odd prime power order $q$ and $V=\mathbb{F}_{q}^{d}$ be the standard $d$-dimensional vector space over $\mathbb{F}_{q}$. Note that all finite vector spaces of odd characteristic are of this form. Equip $V$ with the standard bilinear dot product associated to the nondegenerate quadratic form $Q(v)=v_{1}^{2}+\dots+v_{d}^{2}$. The ”distance” between two points $v$ and $w$ in $V$ as determined by this quadratic form is given by $Q(v-w)$ which is the standard ”Euclidean” distance formula without the square root. The quadratic space $(V,Q)$ will be refered to as $d$-dimensional Euclidean space over $\mathbb{F}_{q}$ in this paper. Note that it is known that there are only two nondegenerate quadratic forms up to isometry on any finite vector space ([16]), besides the one derived from the standard dot product that we are using, the other one is a ”Lorentzian” quadratic form with $Q_{Lorentz}(v)=v_{1}^{2}+\dots+v_{d-1}^{2}+\xi v_{d}^{2}$ where $\xi$ is any fixed nonsquare element in $\mathbb{F}_{q}$. Most of what we discuss will also hold for this other quadratic form but we will stick to the dot product in this paper for brevity. For every $t\in\mathbb{F}_{q}$, we may define the distance-t graph whose vertex set is $V$ and where there is an edge between $v,w\in V$ if and only if $Q(v-w)=t$ and $v\neq w$. This regular graph has a $q^{d}\times q^{d}$ adjacency matrix $\mathbb{A}_{t}$ with respect to some fixed ordering of the vertices of $V$ and common vertex degree $|S_{t}|$ where $S_{t}=\{v\in V\setminus\{0\}|Q(v)=t\}$ is the ”sphere of radius $t$” and has order $q^{d-1}(1+o(1))$ when $t\neq 0$. (The exact order is known but here for brevity we collect secondary terms in $o(1)$ which tends to 00 as $q\to\infty$). In fact this graph is a Cayley graph with connection set $S_{t}$ and hence is in fact vertex-transitive (i.e., there is a graph automorphism taking any vertex to any other vertex). The corresponding Markov chain (see [13] for basic Markov chain terminology) on this graph is called the distance-t random walk in $V$ with the case $t=1$ called the unit random walk in $V$. It corresponds to a situation where at each step, the current state evolves by taking a distance $t$ step with each such step equally likely. The transition matrix for this Markov chain is $T_{t}=\frac{1}{|S_{t}|}\mathbb{A}_{t}$. The matrices $\{\mathbb{A}_{t}|t\in\mathbb{F}_{q}\}$ determine a symmetric association scheme with corresponding real and complex Bose-Messner algebras. What this means is that we have identities of the form
@@
where the intersection numbers $p_{i,j}^{k}$ correspond to the number of ways that a given ”line segment” of length $k$ can be completed to a triangle of side lengths $i,j$ and $k$. We call this particular scheme the Euclidean association scheme for the finite field $\mathbb{F}_{q}$ and recompute certain change of basis matrices $P$ and $Q$ (between a geometric and a spectral basis of the Bose-Messner algebra) which are generally important in the theory of such schemes. We also compute this association scheme’s intersection numbers. (In [12], the higher dimensional numbers were computed in terms of the planar case which we compute explicitly here.) This first part of the paper’s results appear in the first author’s thesis as well as in [12] (some errors in [12] were later corrected in [2] and the connection to Kloosterman sums made more explicit, though these were also already described earlier in [9]). The eigenvalues of $\mathbb{A}_{t}$ are related to twisted Kloosterman sums:
@@
In this paper, we calculate $R_{q,\ell,t}$ (see Theorem 7.1), the probability that you return to the vertex you started from after $\ell$-steps in the random distance $t$-walk, in terms of these moments of Kloosterman sums. We show that this probability is independent of your starting state/vertex and is given by:
@@
Note the first $\frac{1}{q^{2}}$ term in these asymptotic formulas is what one would expect if the location after $\ell$ steps were equally likely to be anywhere in the plane and the second term represents an arithmetic bias against that happening. In particular we see that for $\ell\geq 5$, we have $\mathbb{R}_{q,\ell}$ is $\frac{1}{q^{2}}(1+o(1))$ but for smaller $\ell$ arithmetic bias is significant and so we need at least $5$ steps to achieve uniformity. Note $R_{q,1}=0,R_{q,2}=\frac{1}{q}(1+o(1)),R_{q,4}=\frac{3}{q^{2}}(1+o(1))$ are all ”arithmetically biased”. Similar results hold for $q=1\text{ mod }4$ and can be seen in the paper. Though the main terms in these formulas can be derived using just the spectral gap when $\ell\geq 5$, the explicit nature of the second order term requires, and is equivalent to, the deeper vertical equidistribution results.
@@ 3. Eigenvalues of the Distance $t$-graphs
For any $t\in\mathbb{F}_{q}$, let $X(q,t)$ denote the associated distance $t$-graph on $\mathbb{F}_{q}^{d}$. In this section, we describe the spectrum of the adjacency matrix of this graph, $A_{t}$, when $t\neq\bar{0}$, i.e., for $t\in\mathbb{F}_{q}$. This spectrum was previously computed by various authors for example in [9], [17] but we include a brief discussion here to be self-contained. Standard spectral theory of finite Cayley graphs over Abelian groups, shows that the complex eigenfunctions of this graph’s adjacency operator $A_{t}$ are the characters $\chi_{m}(x)=\chi(m\cdot x)$ as $m$ ranges over $\mathbb{F}_{q}^{d}$. Here $\cdot$ stands for dot product. The corresponding eigenvalue of $A_{t}$ on $\chi_{m}$ which we will denote $\lambda_{m,t,d}$ (recall $d$ is the dimension of $V$) is given by
@@ 5. The P and Q matrices of the Euclidean Association scheme
The calculation of the $P$-matrix of the Euclidean scheme in this section was first done in [12] with corrections in [2] but we provide the details independently here in language more connected to our applications.
@@ 8. Intersection Numbers and Matrices of the scheme
We will concentrate on the planar, $d=2$ case as the higher dimensional cases were reduced to this case in [12]. This case was partially computed in [3] in a different context.